# A preorder on the set of links defined via orbifolds

### v1

@@ 5.2. Small links and determinants
Let $L\subset S^{3}$ be a small link. By Lemma 5.2, the associated $3$-orbifold $\mathcal{O}(L)$ is small. It follows from $\mathit{PSL}(2,\mathbb{C})$-Culler-Shalen theory in the setting of $3$-orbifolds that the $\mathit{PSL}(2,\mathbb{C})$-character variety $\mathcal{X}(\mathcal{O}(L))$ is 0-dimensional, and thus consists of finitely many points. See [22], [15, Sections  3 and 4], and [65, Corollary 10.2].
Suppose, to the contrary, that $\det L=0$. By the proof of Proposition 5.5, $G^{\mathrm{orb}}(L)$ surjects onto the infinite dihedral group $\mathbb{Z}/2\mathbb{Z}\ast\mathbb{Z}/2\mathbb{Z}$. This induces an injection $\mathcal{X}(\mathbb{Z}/2\mathbb{Z}\ast\mathbb{Z}/2\mathbb{Z})\to\mathcal{X}(\mathcal{O}(L))$ between $\mathit{PSL}(2,\mathbb{C})$-character varieties. Here, $\mathbb{Z}/2\mathbb{Z}\ast\mathbb{Z}/2\mathbb{Z}$ surjects onto all finite dihedral groups $D_{2n}$ and those can be realized as subgroups of $\mathit{PSL}(2,\mathbb{C})$ by [46, C.6 and Theorem C.9 in Chapter V] (see also [56, Lemma 4.7]). For $n\geq 2$, the elliptic elements of order $n$ generating the normal subgroup $\mathbb{Z}/n\mathbb{Z}\subset D_{2n}$ have distinct traces. Therefore, $\mathcal{X}(\mathbb{Z}/2\mathbb{Z}\ast\mathbb{Z}/2\mathbb{Z})$ is infinite, which contradicts the fact that $\mathcal{X}(\mathcal{O}(L))$ contains of only finitely many points. ∎
@@
Such a pair $(S^{3},L)$ with $L=L_{1}\sqcup L_{2}$ is a good pair in the sense of Kawauchi. It follows from Kawauchi’s almost identical imitation theory that the pair $(S^{3},L)$ admits almost identical imitations $(S^{3},L^{\ast})$ which are hyperbolic links with the same number of components as $L$ (see [38, Theorem 1.1]). By the construction, the imitation map $q\colon(S^{3},L^{\ast})\to(S^{3},L)$ has the property that the restriction $q|\colon(S^{3},L^{\ast}\setminus{K^{\ast}})\to(S^{3},L\setminus{K})$ is homotopic to a diffeomorphism for any connected components $K^{\ast}$ of $L^{\ast}$ with $q(K^{\ast})=K$. Hence the imitation map induces a proper degree-one map from the exterior $E(L^{\ast})$ onto $E(L)$ which sends the meridians of $L^{\ast}$ to those of $L$, and thus induces an epimorphism $G^{\mathrm{orb}}(L^{\ast})\twoheadrightarrow G^{\mathrm{orb}}(L)$. It follows that $L^{\ast}\succeq L=L_{1}\sqcup L_{2}\succeq L(\frac{p}{q})\sharp L_{2}$ for every $2$-bridge link $L(\frac{p}{q})$.
@@
Let $L\subset S^{3}$ be a small link. By Lemma 5.2, the associated $3$-orbifold $\mathcal{O}(L)$ is small. It follows from $\mathit{PSL}(2,\mathbb{C})$-Culler-Shalen theory [15, Section  4] and [65, Corollary 10.2] that the $\mathit{PSL}(2,\mathbb{C})$-character variety $\mathcal{X}(\mathcal{O}(L))$ is 0-dimensional, and thus consists of finitely many points. By Theorem 1.11(1), any link $L^{\prime}$ such that $L\succeq L^{\prime}$ is small and the orbifold $\mathcal{O}(L^{\prime})$ is small by Lemma 5.2. It follows that the orbifold $\mathcal{O}(L^{\prime})$ is irreducible and atoroidal. Hence, by the geometrization theorem of $3$-orbifolds (see [9, Theorem 2]), $\mathcal{O}(L)$ is hyperbolic or Seifert fibered with a small base which is either hyperbolic, Euclidean, elliptic or bad. We distinguish three mutually exclusive cases according to the geometry of $\mathcal{O}(L^{\prime})$ and prove the finiteness result for each case.
@@
We consider now the case where $\mathcal{O}(L^{\prime})$ is a small Seifert fibered orbifold. Then $L^{\prime}$ is a small Seifert link or a small generalized Montesinos link. A generalized Montesinos link is the union of a classical Montesinos link with $g$ unknotted parallel components which surround the central band as in [11, Figure 7]. For $g\geq 1$, its $2$-fold branched cover is Seifert fibered with a non-orientable base of genus $g$. See [13], [49], [50, Sections 4.7–4.8] for details.
@@
When $\overline{\mathcal{B}^{\prime}}=S^{2}(2,2,n)$, $\Sigma_{2}(L^{\prime})$ is a prism manifold and the order $n$ of the exceptional fiber is bounded above by the order of the first homology group $|H_{1}(\Sigma_{2}(L^{\prime});\mathbb{Z})|=\det L^{\prime}$ which divides $\det L\neq 0$, see [55, Theorem 2(ii), Section 6.2]. Therefore, the orbifold base of $\Sigma_{2}(L^{\prime})$ can take only finitely many topological types, and the proof of Lemma 5.11 shows that $L^{\prime}$ can take only finitely many link types in this case too. If $\overline{\mathcal{B}^{\prime}}=P^{2}(n)$, $\Sigma_{2}(L^{\prime})$ admits also a Seifert fibration with base $S^{2}(2,2,n)$ and thus is a prism manifold, see [55, Theorem 2(vi), Section 6.2]. The finiteness result for $L^{\prime}$ follows from the previous case. ∎
@@ 6. $\pi$-minimal links
By a deep result of Groves, Hull, and Liang [30, Theorem B], this infinite sequence of epimorphisms between the fundamental groups of compact $3$-manifolds eventually contains an isomorphism. Therefore, the sequence of epimorphisms between orbifold groups stabilizes. ∎
@@
If $\frac{1}{\alpha^{\prime}_{1}}+\frac{1}{\alpha^{\prime}_{2}}+\frac{1}{\alpha^{\prime}_{3}}<1$, then $T(\alpha^{\prime}_{1},\alpha^{\prime}_{2},\alpha^{\prime}_{3})\subset\mathit{PSL}(2,\mathbb{R})$ is a hyperbolic triangle group. The two elliptic elements $a=\bar{\varphi}(y)$ and $b=\bar{\varphi}(z)$ of prime orders $\alpha_{2}$ and $\alpha_{3}$ generate the discrete non-abelian group $T(\alpha^{\prime}_{1},\alpha^{\prime}_{2},\alpha^{\prime}_{3})\subset\mathit{PSL}(2,\mathbb{R})$. Then up to taking suitable powers $u=a^{k}$ and $v=b^{\ell}$ to normalize the matrix representatives of $u$ and $v$ in $SL(2,\mathbb{R})$, at least one of Cases (I)–(VII) in [42, Theorem 2.3] holds. Since the triangle group $T(\alpha^{\prime}_{1},\alpha^{\prime}_{2},\alpha^{\prime}_{3})$ is co-compact, Case (II) is impossible by [42, Figures 2 and 3]. Cases (III) and (VI) do not hold because the generators $u$ and $v$ do not have the same order since $\alpha_{2}<\alpha_{3}$. Case (IV) is not possible because none of the generators $u$ or $v$ has order $2$, since $2\leq\alpha_{1}<\alpha_{2}<\alpha_{3}$. If Cases (V) or (VII) happen, then the order of $u$ is $3$ and the order of $v$ is $n\geq 7$ and not divisible by $3$. Hence $\alpha_{2}=3$ and $\alpha_{3}=n$. It follows that $\alpha_{1}=2$ and so $\{\alpha_{1},\alpha_{2},\alpha_{3}\}=\{2,3,n\}$. Moreover, by [42, Section 3, Cases V and VII], we have
@@
We are left with Case (I) where $uv$ has an extreme negative trace. By [42, Proposition 2.2], $u$ and $v$ generate a hyperbolic triangle group $T(\alpha_{2},\alpha_{3},n)$. Therefore, $T(\alpha^{\prime}_{1},\alpha^{\prime}_{2},\alpha^{\prime}_{3})=T(\alpha_{2},\alpha_{3},n)$. The epimorphism $\bar{\varphi}\colon T(\alpha_{1},\alpha_{2},\alpha_{3})\twoheadrightarrow T(\alpha_{2},\alpha_{3},n)$ implies that $\chi(S^{2}(\alpha_{1},\alpha_{2},\alpha_{3}))\geq-\chi(S^{2}(\alpha_{2},\alpha_{3},n))$ by [57, Lemma 2.5]. So, $1-(\frac{1}{\alpha_{1}}+\frac{1}{\alpha_{2}}+\frac{1}{\alpha_{3}})\geq 1-(\frac{1}{\alpha_{2}}+\frac{1}{\alpha_{3}}+\frac{1}{n})$ and hence $n\leq\alpha_{1}$. On the other hand, $\bar{\varphi}(x)$ is an element of prime order $\alpha_{1}$ in $T(\alpha_{2},\alpha_{3},n)$. The only torsion elements in a hyperbolic triangle group are conjugate in a maximal finite cyclic subgroup generated by one of the rotation generators. Therefore, $\alpha_{1}$ must divide one of the numbers $\{\alpha_{2},\alpha_{3},n\}$. Since $\alpha_{2}$ and $\alpha_{3}$ are prime numbers distinct from $\alpha_{1}$, the only possibility is that $\alpha_{1}$ divides $n$. Then $\alpha_{1}=n$ because $n\leq\alpha_{1}$. Therefore, up to permutation, $\{\alpha^{\prime}_{1},\alpha^{\prime}_{2},\alpha^{\prime}_{3}\}=\{\alpha_{1},\alpha_{2},\alpha_{3}\}$ and $\bar{\varphi}\colon T(\alpha_{1},\alpha_{2},\alpha_{3})\twoheadrightarrow T(\alpha^{\prime}_{1},\alpha^{\prime}_{2},\alpha^{\prime}_{3})$ is an isomorphism by the hopfian property of triangle groups. The same argument as in Cases (V) and (VII) shows that $\Sigma_{2}(K)$ and $M$ have isomorphic fundamental groups. ∎
7. Symmetric unions of knots
@@
Let $D$ be an unoriented planar diagram of a knot $K_{D}$ and let $D^{*}$ be the diagram obtained from $D$ by reflecting $D$ across an axis $\Delta$ in the plane. Let $B_{0},B_{1},\dots,B_{k}$ be balls along the axis $\Delta$, each of which is invariant by the reflection and intersects $D$ in a trivial arc. One replaces the trivial tangle $(B_{0},B_{0}\cap(D\cup D^{*}))$ by a $\infty$-tangle to get the connected sum of the diagrams $D$ and $D^{*}$. For $1\leq i\leq k$, one replaces each trivial tangle $(B_{i},B_{i}\cap(D\sharp D^{*}))$ by a $n_{i}$-tangle, where $n_{i}\in\mathbb{Z}$. The knot diagram $(D\cup D^{*})(\infty,n_{1},\dots,n_{k})$ obtained from $D\cup D^{*}$ in this way is called a symmetric union of the diagram $D$ and $D^{*}$. A knot which admits a diagram $(D\cup D^{*})(\infty,n_{1},\dots,n_{k})$ is said to admit a symmetric union presentation with partial knot $K_{D}$, where $K_{D}$ corresponds to the closure of the diagram $D$ such that $(D\cup D^{*})(0,0,\dots,0)=K_{D}\cup K_{D}^{*}$. See Figure 1.
@@
The symmetric union construction always produces ribbon knots and it is still an open question whether there is a ribbon knot that does not admit a symmetric union presentation. One difficulty in finding such an example is that there is no obstruction known for a ribbon knot to be a symmetric union. However, there are candidates like the Montesinos knot $11a_{201}=K(\frac{1}{3},\frac{2}{3},\frac{4}{5})$. In [5, Proposition 1.10], it has been shown that if the knot $11a_{201}$ admits a symmetric union presentation, it has to be skew and the only possible partial knots are $6_{1}$ and $9_{1}$. At the moment, it has not been possible to rule out these two possibilities of partial knots. The goal of this section is to pursue the study of the symmetric union construction by investigating more closely the relationship between the knot types of a symmetric union and of its partial knots.
The following is the key result which relates the $\pi$-orbifold group of a knot and the symmetric union construction. It is due to M. Eisermann (see [43, Theorem 3.3]).
@@
When $K$ is a $2$-bridge knot, $\det K=(\det K_{D})^{2}\neq 1$, then $K_{D}$ is knotted and hence a $2$-bridge knot. The symmetric union is skew because an epimorphism of a $2$-bridge knot group onto the group of a non-trivial knot never kills the longitude by [4, Propositions 1.8 and 1.10].
@@
For a $2$-bridge symmetric union, the finiteness of the possible partial knots has been recently proved by Lamm and Tanaka [44] by using the fact that the partial knots are alternating because they have at most two bridges. See also [62].
Also, as an application of Proposition 7.2 and the orbifold theorem, we obtain the next result, which gives a positive answer to [44, Question (2)].
@@
By [43, Theorem 2.6], $\det K=(\det K_{D})^{2}$ and thus $\det K=\det K_{D}=1$. Since $\varphi_{D}^{\mathrm{orb}}$ kills the image $\bar{\lambda}_{K}$ of the preferred longitude of $K$, $\bar{\lambda}_{K}$ must be trivial in $G^{\mathrm{orb}}(K)$. Then, the contradiction comes from the following lemma. ∎

### v2

@@ 5.2. Small links and determinants
Let $L\subset S^{3}$ be a small link. By Lemma 5.2, the associated $3$-orbifold $\mathcal{O}(L)$ is small. It follows from $\mathit{PSL}(2,\mathbb{C})$-Culler-Shalen theory in the setting of $3$-orbifolds that the $\mathit{PSL}(2,\mathbb{C})$-character variety $\mathcal{X}(\mathcal{O}(L))$ is 0-dimensional, and thus consists of finitely many points. See [22], [15, Sections  3 and 4], and [64, Corollary 10.2].
Suppose, to the contrary, that $\det L=0$. By the proof of Proposition 5.5, $G^{\mathrm{orb}}(L)$ surjects onto the infinite dihedral group $\mathbb{Z}/2\mathbb{Z}\ast\mathbb{Z}/2\mathbb{Z}$. This induces an injection $\mathcal{X}(\mathbb{Z}/2\mathbb{Z}\ast\mathbb{Z}/2\mathbb{Z})\to\mathcal{X}(\mathcal{O}(L))$ between $\mathit{PSL}(2,\mathbb{C})$-character varieties. Here, $\mathbb{Z}/2\mathbb{Z}\ast\mathbb{Z}/2\mathbb{Z}$ surjects onto all finite dihedral groups $D_{2n}$ and those can be realized as subgroups of $\mathit{PSL}(2,\mathbb{C})$ by [45, C.6 and Theorem C.9 in Chapter V] (see also [55, Lemma 4.7]). For $n\geq 2$, the elliptic elements of order $n$ generating the normal subgroup $\mathbb{Z}/n\mathbb{Z}\subset D_{2n}$ have distinct traces. Therefore, $\mathcal{X}(\mathbb{Z}/2\mathbb{Z}\ast\mathbb{Z}/2\mathbb{Z})$ is infinite, which contradicts the fact that $\mathcal{X}(\mathcal{O}(L))$ contains of only finitely many points. ∎
@@
Such a pair $(S^{3},L)$ with $L=L_{1}\sqcup L_{2}$ is a good pair in the sense of Kawauchi. It follows from Kawauchi’s almost identical imitation theory that the pair $(S^{3},L)$ admits almost identical imitations $(S^{3},L^{\ast})$ which are hyperbolic links with the same number of components as $L$ (see [36, Theorem 1.1]). By the construction, the imitation map $q\colon(S^{3},L^{\ast})\to(S^{3},L)$ has the property that the restriction $q|\colon(S^{3},L^{\ast}\setminus{K^{\ast}})\to(S^{3},L\setminus{K})$ is homotopic to a diffeomorphism for any connected components $K^{\ast}$ of $L^{\ast}$ with $q(K^{\ast})=K$. Hence the imitation map induces a proper degree-one map from the exterior $E(L^{\ast})$ onto $E(L)$ which sends the meridians of $L^{\ast}$ to those of $L$, and thus induces an epimorphism $G^{\mathrm{orb}}(L^{\ast})\twoheadrightarrow G^{\mathrm{orb}}(L)$. It follows that $L^{\ast}\succeq L=L_{1}\sqcup L_{2}\succeq L(\frac{p}{q})\sharp L_{2}$ for every $2$-bridge link $L(\frac{p}{q})$.
@@
Let $L\subset S^{3}$ be a small link. By Lemma 5.2, the associated $3$-orbifold $\mathcal{O}(L)$ is small. It follows from $\mathit{PSL}(2,\mathbb{C})$-Culler-Shalen theory [15, Section  4] and [64, Corollary 10.2] that the $\mathit{PSL}(2,\mathbb{C})$-character variety $\mathcal{X}(\mathcal{O}(L))$ is 0-dimensional, and thus consists of finitely many points. By Theorem 1.11(1), any link $L^{\prime}$ such that $L\succeq L^{\prime}$ is small and the orbifold $\mathcal{O}(L^{\prime})$ is small by Lemma 5.2. It follows that the orbifold $\mathcal{O}(L^{\prime})$ is irreducible and atoroidal. Hence, by the geometrization theorem of $3$-orbifolds (see [9, Theorem 2]), $\mathcal{O}(L)$ is hyperbolic or Seifert fibered with a small base which is either hyperbolic, Euclidean, elliptic or bad. We distinguish three mutually exclusive cases according to the geometry of $\mathcal{O}(L^{\prime})$ and prove the finiteness result for each case.
@@
We consider now the case where $\mathcal{O}(L^{\prime})$ is a small Seifert fibered orbifold. Then $L^{\prime}$ is a small Seifert link or a small generalized Montesinos link. A generalized Montesinos link is the union of a classical Montesinos link with $g$ unknotted parallel components which surround the central band as in [11, Figure 7]. For $g\geq 1$, its $2$-fold branched cover is Seifert fibered with a non-orientable base of genus $g$. See [13], [48], [49, Sections 4.7–4.8] for details.
@@
When $\overline{\mathcal{B}^{\prime}}=S^{2}(2,2,n)$, $\Sigma_{2}(L^{\prime})$ is a prism manifold and the order $n$ of the exceptional fiber is bounded above by the order of the first homology group $|H_{1}(\Sigma_{2}(L^{\prime});\mathbb{Z})|=\det L^{\prime}$ which divides $\det L\neq 0$, see [54, Theorem 2(ii), Section 6.2]. Therefore, the orbifold base of $\Sigma_{2}(L^{\prime})$ can take only finitely many topological types, and the proof of Lemma 5.11 shows that $L^{\prime}$ can take only finitely many link types in this case too. If $\overline{\mathcal{B}^{\prime}}=P^{2}(n)$, $\Sigma_{2}(L^{\prime})$ admits also a Seifert fibration with base $S^{2}(2,2,n)$ and thus is a prism manifold, see [54, Theorem 2(vi), Section 6.2]. The finiteness result for $L^{\prime}$ follows from the previous case. ∎
@@ 6. $\pi$-minimal links
By a deep result of Groves, Hull, and Liang [29, Theorem B], this infinite sequence of epimorphisms between the fundamental groups of compact $3$-manifolds eventually contains an isomorphism. Therefore, the sequence of epimorphisms between orbifold groups stabilizes. ∎
@@
If $\frac{1}{\alpha^{\prime}_{1}}+\frac{1}{\alpha^{\prime}_{2}}+\frac{1}{\alpha^{\prime}_{3}}<1$, then $T(\alpha^{\prime}_{1},\alpha^{\prime}_{2},\alpha^{\prime}_{3})\subset\mathit{PSL}(2,\mathbb{R})$ is a hyperbolic triangle group. The two elliptic elements $a=\bar{\varphi}(y)$ and $b=\bar{\varphi}(z)$ of prime orders $\alpha_{2}$ and $\alpha_{3}$ generate the discrete non-abelian group $T(\alpha^{\prime}_{1},\alpha^{\prime}_{2},\alpha^{\prime}_{3})\subset\mathit{PSL}(2,\mathbb{R})$. Then up to taking suitable powers $u=a^{k}$ and $v=b^{\ell}$ to normalize the matrix representatives of $u$ and $v$ in $SL(2,\mathbb{R})$, at least one of Cases (I)–(VII) in [41, Theorem 2.3] holds. Since the triangle group $T(\alpha^{\prime}_{1},\alpha^{\prime}_{2},\alpha^{\prime}_{3})$ is co-compact, Case (II) is impossible by [41, Figures 2 and 3]. Cases (III) and (VI) do not hold because the generators $u$ and $v$ do not have the same order since $\alpha_{2}<\alpha_{3}$. Case (IV) is not possible because neither of the generators $u$ nor $v$ has order $2$, since $2\leq\alpha_{1}<\alpha_{2}<\alpha_{3}$. If Cases (V) or (VII) happen, then the order of $u$ is $3$ and the order of $v$ is $n\geq 7$ and not divisible by $3$. Hence $\alpha_{2}=3$ and $\alpha_{3}=n$. It follows that $\alpha_{1}=2$ and so $\{\alpha_{1},\alpha_{2},\alpha_{3}\}=\{2,3,n\}$. Moreover, by [41, Section 3, Cases V and VII], we have
@@
We are left with Case (I) where $uv$ has an extreme negative trace. By [41, Proposition 2.2], $u$ and $v$ generate a hyperbolic triangle group $T(\alpha_{2},\alpha_{3},n)$. Therefore, $T(\alpha^{\prime}_{1},\alpha^{\prime}_{2},\alpha^{\prime}_{3})=T(\alpha_{2},\alpha_{3},n)$. The epimorphism $\bar{\varphi}\colon T(\alpha_{1},\alpha_{2},\alpha_{3})\twoheadrightarrow T(\alpha_{2},\alpha_{3},n)$ implies that $\chi(S^{2}(\alpha_{1},\alpha_{2},\alpha_{3}))\geq-\chi(S^{2}(\alpha_{2},\alpha_{3},n))$ by [57, Lemma 2.5]. So, $1-(\frac{1}{\alpha_{1}}+\frac{1}{\alpha_{2}}+\frac{1}{\alpha_{3}})\geq 1-(\frac{1}{\alpha_{2}}+\frac{1}{\alpha_{3}}+\frac{1}{n})$ and hence $n\leq\alpha_{1}$. On the other hand, $\bar{\varphi}(x)$ is an element of prime order $\alpha_{1}$ in $T(\alpha_{2},\alpha_{3},n)$. The only torsion elements in a hyperbolic triangle group are conjugate in a maximal finite cyclic subgroup generated by one of the rotation generators. Therefore, $\alpha_{1}$ must divide one of the numbers $\{\alpha_{2},\alpha_{3},n\}$. Since $\alpha_{2}$ and $\alpha_{3}$ are prime numbers distinct from $\alpha_{1}$, the only possibility is that $\alpha_{1}$ divides $n$. Then $\alpha_{1}=n$ because $n\leq\alpha_{1}$. Therefore, up to permutation, $\{\alpha^{\prime}_{1},\alpha^{\prime}_{2},\alpha^{\prime}_{3}\}=\{\alpha_{1},\alpha_{2},\alpha_{3}\}$ and $\bar{\varphi}\colon T(\alpha_{1},\alpha_{2},\alpha_{3})\twoheadrightarrow T(\alpha^{\prime}_{1},\alpha^{\prime}_{2},\alpha^{\prime}_{3})$ is an isomorphism by the hopfian property of triangle groups. The same argument as in Cases (V) and (VII) shows that $\Sigma_{2}(K)$ and $M$ have isomorphic fundamental groups. ∎
7. Applications to symmetric unions of knots
@@ 7.1. Symmetric unions of knots
Let $D$ be an unoriented planar diagram of a knot $K_{D}$ and let $D^{*}$ be the diagram obtained from $D$ by reflecting $D$ across an axis $\Delta$ in the plane. Let $B_{0},B_{1},\dots,B_{k}$ be balls along the axis $\Delta$, each of which is invariant by the reflection and intersects $D$ in a trivial arc. One replaces the trivial tangle $(B_{0},B_{0}\cap(D\cup D^{*}))$ by an $\infty$-tangle to get the connected sum of the diagrams $D$ and $D^{*}$. For $1\leq i\leq k$, one replaces each trivial tangle $(B_{i},B_{i}\cap(D\sharp D^{*}))$ by a $n_{i}$-tangle, where $n_{i}\in\mathbb{Z}$. The knot diagram $(D\cup D^{*})(\infty,n_{1},\dots,n_{k})$ obtained from $D\cup D^{*}$ in this way is called a symmetric union of the diagram $D$ and $D^{*}$. A knot which admits a diagram $(D\cup D^{*})(\infty,n_{1},\dots,n_{k})$ is said to admit a symmetric union presentation with partial knot $K_{D}$, where $K_{D}$ corresponds to the closure of the diagram $D$ such that $(D\cup D^{*})(0,0,\dots,0)=K_{D}\cup K_{D}^{*}$. See Figure 1.
@@
The symmetric union construction always produces ribbon knots, and it is still an open question whether there is a ribbon knot that does not admit a symmetric union presentation. One difficulty in finding such an example is that there is no obstruction known for a ribbon knot to be a symmetric union. However, there are candidates like the Montesinos knot $11a_{201}=K(\frac{1}{3},\frac{2}{3},\frac{4}{5})$. In [5, Proposition 1.10], it has been shown that if the knot $11a_{201}$ admits a symmetric union presentation, it has to be skew and the only possible partial knots are $6_{1}$ and $9_{1}$. At the moment, it has not been possible to rule out these two possibilities of partial knots. The goal of this section is to pursue the study of the symmetric union construction by investigating more closely the relationship between the knot types of a symmetric union and of its partial knots.
The following is the key result which relates the $\pi$-orbifold group of a knot and the symmetric union construction. It is due to M. Eisermann (see [42, Theorem 3.3]).
@@
Note that (1) was already proved in [42, Theorem 3.5]. Regarding (2), $\det K=(\det K_{D})^{2}\neq 1$ implies $K_{D}$ is knotted, and hence a $2$-bridge knot, which can be found in [43, Proposition 3.1]. Furthermore, the symmetric union is skew because an epimorphism of a $2$-bridge knot group onto the group of a non-trivial knot never kills the longitude by [4, Propositions 1.8 and 1.10].
@@
For a $2$-bridge symmetric union, the finiteness of the possible partial knots has been recently proved by Lamm and Tanaka [43] by using the fact that the partial knots are alternating because they have at most two bridges.
7.2. Proof of Theorem 1.12
We now prove Theorem 1.12 which provides obstructions for a $\pi$-minimal knot $K$ to be a symmetric union, in terms of the rank of the homology of its $2$-fold branched cover $\Sigma_{2}(K)$. If $K$ is $\pi$-minimal and admits a symmetric union presentation with partial knot $K_{D}$, then the $\pi$-orbifold group $G^{\mathrm{orb}}(K_{D})$ is isomorphic to either $G^{\mathrm{orb}}(K)$, a dihedral group, or $\mathbb{Z}/2\mathbb{Z}$.
The next proposition shows that $G^{\mathrm{orb}}(K_{D})$ cannot be isomorphic to $G^{\mathrm{orb}}(K)$. It gives a positive answer to [43, Question (2)].
@@
By [42, Theorem 2.6], $\det K=(\det K_{D})^{2}$ and thus $\det K=\det K_{D}=1$. Since $\varphi_{D}^{\mathrm{orb}}$ kills the image $\bar{\lambda}_{K}$ of the preferred longitude of $K$, $\bar{\lambda}_{K}$ must be trivial in $G^{\mathrm{orb}}(K)$. Then, the contradiction comes from the following lemma. ∎